# MedSegMamba: 3D CNN-Mamba Hybrid Architecture for Brain Segmentation

### v2

@@ 1 Introduction
Subcortical brain segmentation is a significant application in medical image processing, as it enables the extraction of quantitative structural information on subcortical regions within MRI scans. This can aid in detecting and tracking morphological deficits in various neuropsychiatric conditions, including Schizophrenia [9], Major Depressive Disorder [14], and Dementia [28]. However, achieving accurate brain segmentation has remained a challenging task due to the intricate 3D structures within the brain, the large number of anatomical labels, and the substantial computational resources required to process scans at full resolution.
@@
While manual segmentation stands as the most trusted method, it is a labor intensive and difficult task, even for experienced clinicians. Automated tools like FreeSurfer [6] have been developed to address these challenges, providing a widely accepted standard for subcortical segmentation. Despite its widespread use, FreeSurfer’s traditional methods can be slow—requiring up to hours to process a single scan—and are often sensitive to data quality issues, which limits their efficiency when dealing with large, heterogeneous datasets.
@@
One of the leading deep learning-based alternatives to FreeSurfer is the FastSurfer pipeline [12] [13], which is capable of performing whole-brain level segmentations. For a 2.5D approach, FastSurfer aggregates three 2D fully convolutional neural networks which utilize the classic encoder-decoder structure originating from the U-Net [24]. While FastSurfer has demonstrated superior performance compared to models such as 3D U-Net [1], QuickNAT [8], and SDNet [25], its reliance on 2D models inherently limits its ability to capture the full 3D spatial dependencies of the brain’s anatomical structures.
@@
On the other hand, 3D patch-based solutions are better suited to capture such geometries. While full 3D volume deep learning models for segmenting many classes are currently not feasible due to data and memory constraints, a patch-based approach is more computationally efficient. It also generates more training samples per subject and better captures local 3D information.
@@
Additionally, traditional pure CNN architectures like used in FastSurfer can suffer from the local receptive fields in each layer, making them more prone to missing out on the full global context in 3D. The Vision Transformer (ViT) architecture [3] addresses this limitation by employing a self-attention mechanism—originally developed for natural language processing [27]—that achieves state-of-the-art performance in image recognition without relying on convolutions. Each self-attention layer has a global receptive field, enabling the model to extract deeper long-range spatial dependencies. For medical image segmentation tasks, hybrid CNN-Transformer architectures modeled after the U-Net have become a popular choice, showing improved generalization and performance [2] [30] [10] [23].
@@
However, the computational cost of Transformers scales quadratically with sequence length, which leads to substantial hardware memory requirements. This makes their application to dense prediction tasks challenging, particularly for large, high-resolution biomedical images. A shifted window-based self-attention approach [17] can mitigate this issue by reducing the computational burden, enabling the use of pure Transformer models. However, it also restricts the receptive fields in each layer, which may result in the loss of global relationships that ViT was designed to capture.
@@
Recent advancements in state space models offer a promising alternative. Mamba [7] introduces the selective scan mechanism (S6) with a hardware aware algorithm to enhance both training and inference efficiency. This allows the model to scale linearly with sequence length while maintaining state-of-the-art performance across various long-sequence processing tasks. Since its introduction, Mamba has been adapted to numerous vision tasks, consistently demonstrating competitive results compared to both Transformer and CNN-based architectures, but with improved memory and parameter efficiency. In 2D applications, Vision Mamba [34] introduces the Vim block, which incorporates a bidirectional State-Space Model (SSM) combined with positional embeddings. Meanwhile, VMamba [16] introduced the Visual State-Space (VSS) block, centered around the 2D-Selective-Scan (SS2D) module. Because the visual components in images are not sequential like language, the Vim block unravels the image in 2 ways (forward and backward) and the SS2D block unravels it along 4 traversal paths, to mitigate any unidirectional bias in the final outputs. MambaAD [11] demonstrates that processing additional traversal paths within each Mamba module can enhance performance in 2D anomaly detection tasks. However, in the realm of 3D segmentation, current methods do not fully exploit the 3D context. For example, U-Mamba [19] unravels the image in only one direction, and SegMamba’s Tri-Oriented Mamba module [32] unfolds the image along just three directions.
@@ 2 Materials and Methods
This study did not require ethical approval as it only utilized publicly available MRI scans previously acquired for studies approved by local institutional review boards, research ethics committees, or human investigation committees.
@@ 2.1 Pipeline
Our pipeline follows a 3D patch-based approach with a hybrid CNN-Mamba model. Initially, the input scans are centered and conformed to LIA orientation, followed by intensity rescaling from 0 to 1, consistent with the steps in the FastSurfer pipeline. These input volumes with dimensions 256 x 256 x 256 are cropped and padded before patch extraction. Patches are then generated with dimensions 96 x 96 x 96, with a step size of 16 voxels between consecutive patches. Each patch is sequentially fed into the model, and the output class probabilities are reconstructed to match the original input image dimensions. The predicted probabilities for each patch are combined through a voting mechanism to determine the class for each voxel, with the final values mapped to the corresponding FreeSurfer labels. This pipeline, as illustrated in Figure 1, ensures that the model can segment an entire scan into 32 classes in under 90 seconds (all 31 subcortical structures covered by FastSurferVINN excluding cortical white matter).
Figure 1: Our pipeline extracts 3D patches from the input scan, feeds them into our model, and reconstructs the output predicted label maps to generate the subcortical segmentation of the input scan.
@@ 2.2 Model Architecture
The selected 3D scanning pattern, depicted in Figure 2, can unravel the volume along 48 unique paths. Our core SS3D module is designed to fully exploit each possible unfolded sequence, with each module processing one of 6 possible groups of 8 sequences.
First, the input volume’s axes are transposed in one of the 6 possible ways (labeled o0 to o5), determined by the assigned orientation index of the SS3D module. The eight sequences are then extracted from the volume by rotating the volume in three different ways, unfolding it along the scanning pattern, and reversing the sequences. These eight sequences are processed independently and in parallel by the S6 blocks, with their outputs reordered and merged to reconstruct the final output volume.
Each method of unraveling the volume (48 in total) is intended to train an independent S6 block to capture the unique geometries best represented by that specific sequence order.
For the S6 blocks, we increase the SSM state dimension from the standard 16 to 64. Although this adjustment slightly reduces processing speed, it enables the model to extract more detailed information about complex structures and possible patch locations, with minimal impact on parameters and memory usage. The SSM feature expansion factor is kept at 1, as the CNN encoder before the bottleneck has already expanded the feature dimension to 1024.
(a) SS3D layout
(b) Scanning pattern
Figure 2: SS3D overview
@@ 2.2.2 VSS3D block
Our 3D Visual State Space (VSS3D) block resembles a traditional Transformer block, with self-attention replaced by our SS3D module, inspired by VMamba. It consists of two residual modules. The first module includes a sequence of layer normalization, linear projection, depth-wise convolution, SiLU activation, SS3D, followed by another layer normalization and linear projection. The second residual module consists of a layer normalization followed by a multilayer perceptron (MLP). This is shown in Figure 3
Figure 3: VSS3D block
@@ 2.2.3 Overall Architecture
The overall model architecture features a 3D CNN encoder and decoder with skip connections, and a series of VSS3D blocks as the bottleneck. Passing through the encoder, four layers of residual blocks and 3 max pooling operations downsample the input patch for an encoded feature tensor. A subsequent convolution brings the channel size to 1024, before the tensor is fed into the bottleneck. The bottleneck comprises a sequence of nine VSS3D blocks, each assigned an orientation index for its inner SS3D module. The output of the bottleneck is normalized and then passed to the decoder, which reconstructs the image to its original input dimensions. Finally, a convolution operation followed by a Softmax activation function are applied to generate a 32-channel output, where each channel represents the probability of an individual class. Each residual block within the encoder and decoder layers consists of a residual connection and two sequences of 3D Convolution, Group Normalization, and Rectified Linear Unit (ReLU). This is visualized in Figure 4.
Figure 4: MedSegMamba model architecture. Only 6 VSS3D layers are shown here, but the model contains 9 in total. The last Conv1x1 and Softmax layers are also not shown.
@@ 2.3 Data
A total of 1784 T1-weighted (T1w) MRI scans were selected from a large-scale heterogeneous dataset representing a uniformly healthy population, compiled from multiple publicly available sources [5]. All scans were acquired at a resolution of 1mm x 1mm x 1mm. Selected subjects came from the Australian Imaging Biomarkers and Lifestyle Study of Ageing (AIBL) [4], Frontotemporal Lobar Degeneration Neuroimaging Initiative (NIFD), Information eXtraction from Images (IXI), Open Access Series of Imaging Studies-1 (OASIS-1) [21], Open Access Series of Imaging Studies-2 (OASIS-2) [20], Southwest University Adult life-span Dataset (SALD) [31], Southwest University Longitudinal Imaging Multimodal Brain Data Repository (SLIM) [15], Parkinson’s Progression Markers Initiative (PPMI) [22], SchizConnect (SchizConnect) [29], and Consortium for Reliability and Reproducibility (CoRR) [35].
Figure 5: Age and Gender Distributions for each Dataset
The dataset was partitioned into training, validation, and test sets with a roughly 3:1:1 ratio. Specifically, the training set included 1076 scans, the validation set contained 345 scans, and the test set comprised 363 scans. A balanced age and gender distribution was achieved across the datasets, as shown in Figure 5. Ground truth segmentations were generated using FreeSurfer, and the input T1w scans were preprocessed with skull-stripping and intensity normalization.
@@ 2.4 Model Training
The described model was trained and evaluated on a 24 GB NVIDIA Quadro 6000 GPU. We employed the AdamW optimizer in conjunction with a Cosine Annealing Warm Restarts learning rate scheduler [18]. The loss function used was a combination of Dice Loss and Weighted Cross Entropy during the first epoch, followed by Dice Loss alone in subsequent epochs. Training was conducted on 27 patches per scan (three steps per dimension), utilizing gradient accumulation to simulate each image as a single batch, accounting for the variability in class presence across patches. A dropout rate of 0.1 and a drop path rate of 0.3 were applied within the bottleneck.
@@ 2.5 Model Evaluation and Statistical Test
To evaluate the efficacy of our novel SS3D module over other 3D mamba-based vision modules for latent space learning, we also trained a model identical to our MedSegMamba architecture, except substituting the first residual module in each VSS3D block with SegMamba’s Tri-oriented Mamba module. The resulting block is identical to SegMamba’s Tri-oriented Spatial Mamba block except excluding the Gated Spatial Convolution module, in order to focus specifically on evaluating the core Mamba modules. This model, referred to as SegMambaBot, was trained following the same procedure as MedSegMamba.
MedSegMamba was also evaluated against a retrained TABSurfer and pretrained FastSurferVINN (obtained from the FastSurfer GitHub repository) using 363 FreeSurfer segmentations as ground truths. Both SegMambaBot and TABSurfer were trained using 3D patch-based approaches with input patch sizes identical to those used in MedSegMamba.
The evaluation metrics included the Dice Similarity Coefficient (DSC), Volume Similarity (VS) [26], and Average Symmetric Surface Distance (ASSD) [33], assessing both the overall similarity of the segmentations and the contour quality relative to the ground truth. The Wilcoxon signed-rank test was employed to assess the significance of improvements in MedSegMamba’s performance over the other models for each metric. Statistical significance was set to P < 0.05.
@@ 3 Results
Figure 6: 3D renderings for a sample segmented by each method.
The average metrics from evaluating MedSegMamba, SegMambaBot, TABSurfer, and FastSurferVINN against the FreeSurfer-generated ground truths are displayed in Table 1. MedSegMamba consistently achieved high scores in Dice Similarity Coefficient (DSC), Volume Similarity (VS), and Average Symmetric Surface Distance (ASSD) across all datasets, demonstrating superior overall performance. MedSegMamba’s improvement was statistically significant across all metrics, except for the DSC and VS metrics when compared to SegMamba (P-values of 0.114 and 0.425, respectively).

### v3

@@ 1 Introduction
Subcortical brain segmentation is a significant application in medical image processing, as it enables the extraction of quantitative structural information from MRI scans. These localized details can aid in detecting and monitoring morphological deficits in various neuropsychiatric conditions, including Schizophrenia [10], Major Depressive Disorder [15], and Dementia [30]. However, achieving accurate brain segmentation has remained a formidable challenge due to the intricate 3D structures within the brain, the large number of anatomical labels, and the substantial computational resources required to process MRI scans at full resolution.
@@
While manual segmentation stands as the most trusted method, it is a tedious and difficult task, even for experienced clinicians. Automated tools like FreeSurfer [7] have been developed to address these challenges, providing a widely accepted standard for subcortical segmentation. Despite its widespread use, FreeSurfer’s traditional methods can be slow—requiring many hours to process a single scan—and are often sensitive to data quality issues, which limits their reliability when processing large, heterogeneous datasets.
@@
The FastSurfer pipeline [13] [14] is recognized as one of the foremost deep learning-based alternatives to FreeSurfer, capable of executing whole-brain level segmentations. For a 2.5D approach, FastSurfer aggregates three 2D fully convolutional neural networks that utilize the classic encoder-decoder structure originating from the U-Net [26]. Even though FastSurfer has demonstrated superior performance compared to standard alternative models such as 3D U-Net [2], QuickNAT [9], and SDNet [27], its reliance on 2D models inherently limits its ability to capture the full 3D spatial dependencies of the brain’s anatomical structures.
@@
On the other hand, 3D patch-based solutions are better suited to capture such geometries. Although full 3D volume deep learning models for segmenting many classes are currently not feasible due to data and memory constraints, a patch-based approach significantly reduces memory usage and generates more training samples per subject. Additionally, smaller patches allow the model to better capture local 3D information, leading to more accurate segmentation in complex anatomical structures.
@@
Moreover, traditional pure CNN architectures like those used in FastSurfer can suffer from the local receptive fields in each layer, making them susceptible to overlooking the comprehensive global 3D context. The Vision Transformer (ViT) architecture [4] addresses this limitation by employing a self-attention mechanism—originally developed for natural language processing [29]—that achieves state-of-the-art performance in image recognition without relying on convolutions. Each self-attention layer has a global receptive field, enabling the model to extract deeper long-range spatial dependencies. Hybrid CNN-Transformer architectures modeled after the U-Net have become popular for medical image segmentation tasks, showing improved generalization and performance [3] [11]. Specifically for subcortical segmentation, TABSurfer [1], a CNN-Transformer hybrid building off of TABS [25] and inspired by TransBTS [32], demonstrated the advantages of combining a hybrid architecture with a 3D patch-based approach.
@@
However, the computational cost of Transformers scales quadratically with sequence length, which leads to substantial hardware memory requirements. This makes their application to dense prediction tasks challenging, particularly for large, high-resolution biomedical images. A shifted window-based self-attention approach [19] can mitigate this issue by reducing the computational burden, enabling the use of pure Transformer models. However, it also restricts the receptive fields in each layer, potentially leading to the loss of global relationships that ViT was designed to capture.
Recent advancements in state space models offer a promising alternative. Mamba [8] introduces the selective scan mechanism (S6) with a hardware-aware algorithm to enhance both training and inference efficiency. This innovation allows the model to scale linearly with respect to sequence length while maintaining state-of-the-art performance across various long-sequence processing tasks.
@@
Since its introduction, Mamba has been adapted to various vision tasks, consistently demonstrating competitive results compared to both Transformer and CNN-based architectures, but with improved memory and parameter efficiency. For 2D image applications, Vision Mamba [36] introduces the Vim block, which incorporates a bidirectional State-Space Model (SSM) combined with positional embeddings. Meanwhile, VMamba [18] introduced the Visual State-Space (VSS) block, centered around the 2D-Selective-Scan (SS2D) module. Because the visual components in images are not sequential, unlike the ordered nature of text and audio, the Vim block unravels the image in two ways (forward and backward) and the SS2D module unravels the image along four traversal paths, to mitigate any unidirectional bias in the final outputs. MambaAD [12] demonstrates that processing additional traversal paths within each Mamba module can enhance performance in 2D anomaly detection tasks. However, in the realm of 3D segmentation, current methods do not fully exploit the richness of the 3D context. For example, U-Mamba [21] unravels the image in only one direction, and SegMamba’s Tri-Oriented Mamba module [34] unfolds the image along just three directions.
@@ 2 Materials and Methods
This study did not require ethical approval as it utilized publicly available anonymous MRI scans previously acquired for studies approved by local institutional review boards, research ethics committees, or human investigation committees.
@@ 2.1 Pipeline
The subcortical segmentation pipeline employs a 3D patch-based approach with a hybrid CNN-Mamba model. Initially, the input scans are centered and conformed to LIA orientation, followed by intensity rescaling from 0 to 1, consistent with the steps in the FastSurfer pipeline. These input volumes with dimensions 256 x 256 x 256 are cropped and padded before patch extraction. Patches are then extracted with dimensions 96 x 96 x 96, with a step size of 16 voxels between consecutive patches. Each patch is sequentially fed into the model, and the output class probabilities are reconstructed to match the original input image dimensions. The predicted probabilities for each patch are combined through a voting mechanism to determine the class for each voxel, with the final values mapped to the corresponding FreeSurfer labels. As illustrated in Figure 1a, this pipeline ensures that the model can segment an entire scan into 32 classes in under 90 seconds (all 31 subcortical structures covered by FastSurferVINN excluding cortical white matter).
The hippocampal subfield segmentation process is visualized in Figure 1b. Using the left and right hippocampus classes in the subcortical segmentation, a bounding box with dimensions 96 x 96 x 96 is drawn around the hippocampus in the input scan, and this patch is fed to a separate hybrid CNN-Mamba model, which segments the hippocampal subfields. The output is then padded to the original input image dimensions. This segments the hippocampus into 25 classes according to FreeSurfer’s “FS60” level of hierarchy except with left and right labels separated.
Figure 1: (a) The subcortical segmentation pipeline extracts 3D patches from the input scan, feeds them into the model, and reconstructs the output predicted label maps to generate the subcortical segmentation of the input scan. (b) The hippocampal subfield segmentation pipeline uses the subcortical segmentation label map to extract one patch centered on the hippocampus region. This patch is fed into the model and the output is padded to the original shape.
@@ 2.2 Model Architecture
Figure 2: (a) The SS3D module encodes 8 sequences from the input volume, processes each with an independent S6 block, and then merges the outputs. (b) VSS3D block layout, consisting of SS3D and MLP residual modules. (c) The MedSegMamba model architecture has an encoder-decoder structure with a hybrid of CNN and Mamba-based blocks. Only 6 VSS3D layers are shown here, but the model contains 9 in total. The last Conv1x1 and Softmax layers are also not shown. (d) The SS3D modules unravel the sequences along this continuous 3D scanning pattern.
@@
The selected 3D scanning pattern, depicted in Figure 2d, can unravel the volume along 48 unique traversal paths. Our core SS3D module, shown in Figure 2a, is designed to fully exploit each possible unfolded sequence, with each module processing one of six possible groups of eight sequences.
First, the input volume’s axes are transposed in one of the six possible ways (labeled o0 to o5), determined by the assigned orientation index of the SS3D module. The eight sequences are then extracted from the volume by rotating the volume in three different ways, unfolding it along the scanning pattern, and reversing the sequences. These eight sequences are processed independently and in parallel by the S6 blocks, with their outputs reordered and merged to reconstruct the final output volume. Each of the 48 methods of unraveling the volume is intended to train independent S6 blocks to capture the unique geometries best represented by that specific sequence order.
For the S6 blocks, we increase the SSM state dimension from the standard 16 to 64. Although this adjustment slightly reduces processing speed, it enables the model to extract more detailed information about complex structures with minimal impact on parameters and memory usage. The SSM feature expansion factor is set to 1 to keep the dimensionality constant, as the CNN encoder before the bottleneck has already expanded the feature dimension to 1024.
@@ 2.2.2 VSS3D block
Our 3D Visual State Space (VSS3D) block resembles a traditional Transformer block, with self-attention replaced by our SS3D module, inspired by VMamba. It consists of two residual modules. The first module includes a sequence of layer normalization, linear projection, depth-wise convolution, SiLU activation, and SS3D, followed by another layer normalization and linear projection. The second residual module consists of a layer normalization followed by a multilayer perceptron (MLP). This is shown in Figure 2b.
@@ 2.2.3 Overall Architecture
The overall model architecture features a 3D CNN encoder and decoder with skip connections and a series of VSS3D blocks as the bottleneck. Passing through the encoder, four layers of residual blocks and 3 max pooling operations downsample the input patch for an encoded feature tensor. A subsequent convolution brings the channel size to 1024, before the tensor is fed into the bottleneck. The bottleneck comprises a sequence of nine VSS3D blocks, each assigned an orientation index for its inner SS3D module. The output of the bottleneck is normalized and then passed to the decoder, which reconstructs the image to its original input dimensions. Finally, a convolution operation and a Softmax activation function are applied to generate either a 25 or 32 channel output, where each channel represents the probability of an individual class. The only difference between the subcortical and hippocampal subfield segmentation models is in the channel sizes of these last two layers. Each residual block within the encoder and decoder layers consists of a residual connection and two sequences of 3D Convolution, Group Normalization, and Rectified Linear Unit (ReLU). This is visualized in Figure 2c.
@@ 2.3 Data
For the subcortical segmentation dataset, a total of 1784 T1-weighted (T1w) MRI scans were selected from a large-scale heterogeneous dataset representing a uniformly healthy population, compiled from multiple publicly available sources [6]. All scans were acquired at a resolution of 1mm x 1mm x 1mm. Selected subjects came from the Australian Imaging Biomarkers and Lifestyle Study of Ageing (AIBL) [5], Frontotemporal Lobar Degeneration Neuroimaging Initiative (NIFD) database (http://4rtniftldni.ini.usc.edu), Information eXtraction from Images (IXI) (http://brain-development.org/ixi-dataset), Open Access Series of Imaging Studies-1 (OASIS-1) [23], Open Access Series of Imaging Studies-2 (OASIS-2) [22], Southwest University Adult life-span Dataset (SALD) [33], Southwest University Longitudinal Imaging Multimodal Brain Data Repository (SLIM) [17], Parkinson’s Progression Markers Initiative (PPMI) [24], SchizConnect (SchizConnect) [31], and Consortium for Reliability and Reproducibility (CoRR) [37].
For the hippocampal subfield segmentation dataset, 1221 1mm x 1mm x 1mm T1w scans were selected from the same large-scale multi-site dataset [6]. These subjects came from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu) and Brain Genomics Superstruct Project (BGSP) [16], in addition to AIBL, IXI, NIFD, and OAS1.
Both datasets were partitioned into training, validation, and test sets with a roughly 3:1:1 ratio. Specifically, the subcortical dataset had a 1076/345/363 split, and the hippocampus dataset had a 719/254/248 split. A roughly balanced age and gender distribution was achieved across the datasets, as shown in Figure 3.
Ground truth segmentations were generated using FreeSurfer, and the input T1w scans were preprocessed with skull-stripping and intensity normalization.
Figure 3: (a) Age and Gender Distributions for the Subcortical Segmentation Dataset (b) Age and Gender Distributions for the Hippocampal Subfield Segmentation Dataset.
@@ 2.4 Model Training
The described model was trained and evaluated on a 24 GB NVIDIA Quadro 6000 GPU. We employed the AdamW optimizer in conjunction with a Cosine Annealing Warm Restarts learning rate scheduler [20]. The loss function used was a combination of Dice Loss and Weighted Cross Entropy during the first epoch, followed by Dice Loss alone in subsequent epochs. Subcortical segmentation training was conducted for 15 epochs on 27 patches per scan (three steps per dimension), utilizing gradient accumulation to simulate each image as a single batch, accounting for the variability in class presence across patches. Hippocampal subfield segmentation training followed the same procedure for 63 epochs except with a batch size of 1. A dropout rate of 0.1 and a drop path rate of 0.3 were applied within the bottleneck.
@@ 2.5 Model Evaluation and Statistical Test
To evaluate the efficacy of our novel SS3D module over other 3D mamba-based vision modules for latent space learning, we also trained a model identical to our MedSegMamba architecture, except substituting the first residual module in each VSS3D block with SegMamba’s Tri-oriented Mamba module. The resulting block is identical to SegMamba’s Tri-oriented Spatial Mamba block except without the Gated Spatial Convolution module, in order to focus specifically on evaluating the core Mamba modules. This model, referred to as SegMambaBot, was trained following the same procedures as MedSegMamba.
For subcortical segmentation, MedSegMamba was also evaluated against a pretrained FastSurferVINN (obtained from the FastSurfer GitHub repository) and a retrained TABSurfer. Both SegMambaBot and TABSurfer were trained using 3D patch-based approaches with input patch sizes identical to those used in MedSegMamba.
For hippocampal subfield segmentation, MedSegMamba was evaluated against SegMambaBot and TABSurfer, with each model being identical to their subcortical segmentation counterparts except for the final layers generating a 25 instead of 32 channel output. Each model was trained following the same procedure. FastSurfer was not evaluated for this task because only one patch containing the hippocampus needs to be processed, which is small enough for a 3D model to handle in one pass. Aggregating multiple 2D slice-based networks to process this smaller region would be slower and unnecessary.
The evaluation metrics included the Dice Similarity Coefficient (DSC), Volume Similarity (VS) [28], and Average Symmetric Surface Distance (ASSD) [35], assessing both the overall similarity of the segmentations and the contour quality relative to the ground truth. The Wilcoxon signed-rank test was employed to evaluate the significance of improvements in MedSegMamba’s performance over the other models for each metric. Statistical significance was set to P < 0.05.
@@ 3 Results
3.1 Subcortical Segmentation
Figure 4: 2D slices of a sample’s subcortical segmentation by each method.
The average metrics used for evaluating MedSegMamba, SegMambaBot, TABSurfer, and FastSurferVINN against the FreeSurfer-generated subcortical segmentation ground truths are displayed in Table 1. MedSegMamba consistently achieved high scores in Dice Similarity Coefficient (DSC), Volume Similarity (VS), and Average Symmetric Surface Distance (ASSD) across all datasets, demonstrating superior overall performance. MedSegMamba’s improvement was statistically significant across all metrics, except for the DSC and VS metrics when compared to SegMamba (P-values of 0.114 and 0.425, respectively).